OpenStudy (anonymous):

PLEASE help!!! medal awarded!! How can x + 8 = -x + 7 be set up as a system of equations? a.) y = x + 8 y = -x + 7 b.) y = x + 8 y = -3x + 7 c.) y = x + 5 y = -2x + 11 d.) y = 8 y = -3x + 7

OpenStudy (the_fizicx99):

A ;p

OpenStudy (anonymous):

how do you know

OpenStudy (the_fizicx99):

Because, if y = x + 8 and y = -x + 7 Then plug in x + 8 for y into y = -x + 7 x+8 = -x + 7

OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 15x - y = -6 5x - 3y = -13

OpenStudy (the_fizicx99):

Make one question per post, it lags up :c